OpenStudy (anonymous):

At the movie theatre, child admission is$5.40 and adult admission is $9.90 . On Tuesday, 131 tickets were sold for a total sales of 954.90 . How many child tickets were sold that day?

Directrix (directrix):

Let C represent the number of children tickets that were puchased Let A represent the number of adult tickets that were purchased. So, (1) C + A = 131 The revenue from child tickets is 5.4*C The revenue from adult tickets is 9.9 * A So, (2) 5.4*C + 9.9*A = 954.9

Directrix (directrix):

The task is to solve these two simultaneous equations (1) C + A = 131 (2) 5.4*C + 9.9*A = 954.9 ---------------------------- @clairelynn93

Directrix (directrix):

C + A = 131 C = 131 - A 5.4(131 - A) + 9.9*A = 954.9 707.4 - 5.4 A + 9.9 A = 954.9 Solve the above equation for A to get the number of adults.